OpenStudy (anonymous):

Solve the following system of equations 2x + 5y = −3 −15x − 5y = −10 Show all your work.

OpenStudy (anonymous):

Omg someone help me please

OpenStudy (anonymous):

just add down. so 2x + (-15x) = 5y+(-5y) = (-3)+(-10) = then solve for x, plug in, and solve for y

OpenStudy (anonymous):

The easiest way to solve this is by elimination. 2x + 5y = −3 −15x − 5y = −10 Add vertically. 2x + 5y = −3 −15x − 5y = −10 ---------------- -13x = -13 x = 1 Now, plug this back in and solve for y. 2x + 5y = −3 2(1) + 5y = −3 Solve for y.

OpenStudy (anonymous):

can some one please just give me the whole answer. Pleaseeeeeee

OpenStudy (anonymous):

Do you want a step-by-step answer for the process?

OpenStudy (anonymous):

you... don't know how to plug in x=1 into either equation to solve for y...?

OpenStudy (anonymous):

@rerite2 yes pleaseeeee

OpenStudy (anonymous):

I will do my best to give you a breakdown of the thought process required to do this. It may seem like I'm being very simplistic but I think that's what you want. Ready?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

The easiest way to solve this is by elimination. 2x + 5y = −3 −15x − 5y = −10 Add vertically. 2x + 5y = −3 −15x − 5y = −10 ---------------- -13x = -13 x = 1 Now, plug this back in and solve for y. 2x + 5y = −3 2(1) + 5y = −3 5y = -3-2 5y = -5 y = -1 and please check what dyslexic means..

OpenStudy (anonymous):

Okay. First, a few questions. You know that an equation has an equal sign and what is done on one side of the equal sign must be done on the other side, right? So each side is balanced, i.e., equal. Understand?